OpenStudy (anonymous):

f(x)=1-6x solve f(x)=0

OpenStudy (aroub):

1

OpenStudy (anonymous):

plug in o for x. so 1-6(0)=1

jimthompson5910 (jim_thompson5910):

f(x)=1-6x 0=1-6x -1=-6x -1/(-6)=x 1/6=x x=1/6 So the solution is \[\large x=\frac{1}{6}\]

OpenStudy (anonymous):

but its a function in terms of x

jimthompson5910 (jim_thompson5910):

exactly, but they want to know which values of x make f(x) equal to zero not f(0)

OpenStudy (anonymous):

oh i see